OpenStudy (anonymous):

Find conditions on k that will make the following system of equations have a unique solution. 3kx+2y = 2 12x+2ky = 4 What is the unique soln of x and y?

OpenStudy (arbershabani97):

answer is k=2 and x=1 and y=-2 but still I can't understand it, I just tested numbers, sorry for telling just the answer, but it may help find the way to find the solution its interesting question

OpenStudy (anonymous):

well, lets work on it together! haha my first inital thought would be to put it into a matrix. however, my only limitation would be that k in the matrix

OpenStudy (anonymous):

in this instances, cramer's rule

OpenStudy (arbershabani97):

i don't know how can we write K in the matrix

OpenStudy (anonymous):

\[\left[\begin{matrix}3k &2 \\ 12 & 2k\end{matrix}\right] = \left[\begin{matrix}4 \\ 2\end{matrix}\right]\]

OpenStudy (anonymous):

but yeah, i dunno how to go from there

OpenStudy (arbershabani97):

i tried to solve it, and I probably can but it will be so complicated that my head is not ready to survive it :P ahahahha sorry

OpenStudy (anonymous):

hahaha. it is all good. thanks tho. ill give u a medal for trying :)

OpenStudy (arbershabani97):

thanks man :)

OpenStudy (anonymous):

ill keep trying to solve it